OpenStudy (anonymous):

Which of these can be used to find forty-four percent of eighty-two? Written as equivalent ratios, forty-four over one hundred equals x over eighty-two. Written as equivalent ratios, forty-four over eighty-two equals x over one hundred. Written as equivalent ratios, forty-four over one hundred equals eighty-two over x. Written as equivalent ratios, one hundred over eighty-two equals x over forty-four.

OpenStudy (anonymous):

.44 x 82 = 36.08 or 44/100 = x/82 100x = (44)(82) 100x = 3608 x = 3608/100 x = 36.08 answer : A

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

your quite welcome :)